\begin{thebibliography}{41}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Abdul{-}Mageed et~al.(2021)Abdul{-}Mageed, Elmadany, and
  Nagoudi}]{Abdul-MageedEN20}
Muhammad Abdul{-}Mageed, AbdelRahim~A. Elmadany, and El~Moatez~Billah Nagoudi.
  2021.
\newblock \href {https://doi.org/10.18653/V1/2021.ACL-LONG.551} {{ARBERT} {\&}
  {MARBERT:} deep bidirectional transformers for arabic}.
\newblock In \emph{Proceedings of the 59th Annual Meeting of the Association
  for Computational Linguistics and the 11th International Joint Conference on
  Natural Language Processing, {ACL/IJCNLP} 2021, (Volume 1: Long Papers),
  Virtual Event, August 1-6, 2021}, pages 7088--7105. Association for
  Computational Linguistics.

\bibitem[{Abdul-Mageed et~al.(2020)Abdul-Mageed, Zhang, Bouamor, and
  Habash}]{abdul-mageed-etal-2020-nadi}
Muhammad Abdul-Mageed, Chiyu Zhang, Houda Bouamor, and Nizar Habash. 2020.
\newblock \href {https://aclanthology.org/2020.wanlp-1.9} {{NADI} 2020: The
  first nuanced {A}rabic dialect identification shared task}.
\newblock In \emph{Proceedings of the Fifth Arabic Natural Language Processing
  Workshop}, pages 97--110, Barcelona, Spain (Online). Association for
  Computational Linguistics.

\bibitem[{Adiwardana et~al.(2020)Adiwardana, Luong, So, Hall, Fiedel,
  Thoppilan, Yang, Kulshreshtha, Nemade, Lu, and Le}]{abs-2001-09977}
Daniel Adiwardana, Minh{-}Thang Luong, David~R. So, Jamie Hall, Noah Fiedel,
  Romal Thoppilan, Zi~Yang, Apoorv Kulshreshtha, Gaurav Nemade, Yifeng Lu, and
  Quoc~V. Le. 2020.
\newblock \href {http://arxiv.org/abs/2001.09977} {Towards a human-like
  open-domain chatbot}.
\newblock \emph{CoRR}, abs/2001.09977.

\bibitem[{Ahmed et~al.(2023)Ahmed, Pan, Bo, Su, Cao, Zhang, and Liu}]{rnt}
Murtadha Ahmed, Shengfeng Pan, Wen Bo, Jianlin Su, Xinxin Cao, Wenze Zhang, and
  Yunfeng Liu. 2023.
\newblock \href {https://doi.org/10.1162/TACL\_A\_00574} {Rank-aware negative
  training for semi-supervised text classification}.
\newblock \emph{Trans. Assoc. Comput. Linguistics}, 11:771--786.

\bibitem[{Alharbi et~al.(2020)Alharbi, Alamro, Alshehri, Khayyat, Kalkatawi,
  Jaber, and Zhang}]{abs-2011-00578}
Basma Alharbi, Hind Alamro, Manal Alshehri, Zuhair Khayyat, Manal Kalkatawi,
  Inji~Ibrahim Jaber, and Xiangliang Zhang. 2020.
\newblock \href {http://arxiv.org/abs/2011.00578} {{ASAD:} {A} twitter-based
  benchmark arabic sentiment analysis dataset}.
\newblock \emph{CoRR}, abs/2011.00578.

\bibitem[{Aly and Atiya(2013)}]{AlyA13}
Mohamed~A. Aly and Amir~F. Atiya. 2013.
\newblock \href {https://aclanthology.org/P13-2088/} {{LABR:} {A} large scale
  arabic book reviews dataset}.
\newblock In \emph{Proceedings of the 51st Annual Meeting of the Association
  for Computational Linguistics, {ACL} 2013, 4-9 August 2013, Sofia, Bulgaria,
  Volume 2: Short Papers}, pages 494--498. The Association for Computer
  Linguistics.

\bibitem[{Antoun et~al.(2020)Antoun, Baly, and Hajj}]{antoun2020arabert}
Wissam Antoun, Fady Baly, and Hazem Hajj. 2020.
\newblock \href {https://aclanthology.org/2020.osact-1.2} {{A}ra{BERT}:
  Transformer-based model for {A}rabic language understanding}.
\newblock In \emph{Proceedings of the 4th Workshop on Open-Source Arabic
  Corpora and Processing Tools, with a Shared Task on Offensive Language
  Detection}, pages 9--15, Marseille, France. European Language Resource
  Association.

\bibitem[{Bouamor et~al.(2019)Bouamor, Hassan, and Habash}]{BouamorHH19}
Houda Bouamor, Sabit Hassan, and Nizar Habash. 2019.
\newblock \href {https://doi.org/10.18653/V1/W19-4622} {The {MADAR} shared task
  on arabic fine-grained dialect identification}.
\newblock In \emph{Proceedings of the Fourth Arabic Natural Language Processing
  Workshop, WANLP@ACL 2019, Florence, Italy, August 1, 2019}, pages 199--207.
  Association for Computational Linguistics.

\bibitem[{Brown et~al.(2020)Brown, Mann, Ryder, Subbiah, Kaplan, Dhariwal,
  Neelakantan, Shyam, Sastry, Askell, Agarwal, Herbert{-}Voss, Krueger,
  Henighan, Child, Ramesh, Ziegler, Wu, Winter, Hesse, Chen, Sigler, Litwin,
  Gray, Chess, Clark, Berner, McCandlish, Radford, Sutskever, and
  Amodei}]{gpt3}
Tom~B. Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared Kaplan,
  Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda
  Askell, Sandhini Agarwal, Ariel Herbert{-}Voss, Gretchen Krueger, Tom
  Henighan, Rewon Child, Aditya Ramesh, Daniel~M. Ziegler, Jeffrey Wu, Clemens
  Winter, Christopher Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott
  Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec
  Radford, Ilya Sutskever, and Dario Amodei. 2020.
\newblock \href
  {https://proceedings.neurips.cc/paper/2020/hash/1457c0d6bfcb4967418bfb8ac142f64a-Abstract.html}
  {Language models are few-shot learners}.
\newblock In \emph{Advances in Neural Information Processing Systems 33: Annual
  Conference on Neural Information Processing Systems 2020, NeurIPS 2020,
  December 6-12, 2020, virtual}.

\bibitem[{Chowdhery et~al.(2023)Chowdhery, Narang, Devlin, Bosma, Mishra,
  Roberts, Barham, Chung, Sutton, Gehrmann, Schuh, Shi, Tsvyashchenko, Maynez,
  Rao, Barnes, Tay, Shazeer, Prabhakaran, Reif, Du, Hutchinson, Pope, Bradbury,
  Austin, Isard, Gur{-}Ari, Yin, Duke, Levskaya, Ghemawat, Dev, Michalewski,
  Garcia, Misra, Robinson, Fedus, Zhou, Ippolito, Luan, Lim, Zoph, Spiridonov,
  Sepassi, Dohan, Agrawal, Omernick, Dai, Pillai, Pellat, Lewkowycz, Moreira,
  Child, Polozov, Lee, Zhou, Wang, Saeta, Diaz, Firat, Catasta, Wei,
  Meier{-}Hellstern, Eck, Dean, Petrov, and Fiedel}]{palm}
Aakanksha Chowdhery, Sharan Narang, Jacob Devlin, Maarten Bosma, Gaurav Mishra,
  Adam Roberts, Paul Barham, Hyung~Won Chung, Charles Sutton, Sebastian
  Gehrmann, Parker Schuh, Kensen Shi, Sasha Tsvyashchenko, Joshua Maynez,
  Abhishek Rao, Parker Barnes, Yi~Tay, Noam Shazeer, Vinodkumar Prabhakaran,
  Emily Reif, Nan Du, Ben Hutchinson, Reiner Pope, James Bradbury, Jacob
  Austin, Michael Isard, Guy Gur{-}Ari, Pengcheng Yin, Toju Duke, Anselm
  Levskaya, Sanjay Ghemawat, Sunipa Dev, Henryk Michalewski, Xavier Garcia,
  Vedant Misra, Kevin Robinson, Liam Fedus, Denny Zhou, Daphne Ippolito, David
  Luan, Hyeontaek Lim, Barret Zoph, Alexander Spiridonov, Ryan Sepassi, David
  Dohan, Shivani Agrawal, Mark Omernick, Andrew~M. Dai,
  Thanumalayan~Sankaranarayana Pillai, Marie Pellat, Aitor Lewkowycz, Erica
  Moreira, Rewon Child, Oleksandr Polozov, Katherine Lee, Zongwei Zhou, Xuezhi
  Wang, Brennan Saeta, Mark Diaz, Orhan Firat, Michele Catasta, Jason Wei,
  Kathy Meier{-}Hellstern, Douglas Eck, Jeff Dean, Slav Petrov, and Noah
  Fiedel. 2023.
\newblock \href {http://jmlr.org/papers/v24/22-1144.html} {Palm: Scaling
  language modeling with pathways}.
\newblock \emph{J. Mach. Learn. Res.}, 24:240:1--240:113.

\bibitem[{Conneau et~al.(2020)Conneau, Khandelwal, Goyal, Chaudhary, Wenzek,
  Guzm{\'a}n, Grave, Ott, Zettlemoyer, and
  Stoyanov}]{conneau-etal-2020-unsupervised}
Alexis Conneau, Kartikay Khandelwal, Naman Goyal, Vishrav Chaudhary, Guillaume
  Wenzek, Francisco Guzm{\'a}n, Edouard Grave, Myle Ott, Luke Zettlemoyer, and
  Veselin Stoyanov. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.747} {{Unsupervised
  Cross-lingual Representation Learning at Scale}}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 8440--8451, Online. Association for
  Computational Linguistics.

\bibitem[{Dadas et~al.(2020)Dadas, Perelkiewicz, and Poswiata}]{DadasPP20}
Slawomir Dadas, Michal Perelkiewicz, and Rafal Poswiata. 2020.
\newblock \href {https://doi.org/10.1007/978-3-030-61534-5\_27} {Pre-training
  polish transformer-based language models at scale}.
\newblock In \emph{Artificial Intelligence and Soft Computing - 19th
  International Conference, {ICAISC} 2020, Zakopane, Poland, October 12-14,
  2020, Proceedings, Part {II}}, volume 12416 of \emph{Lecture Notes in
  Computer Science}, pages 301--314. Springer.

\bibitem[{de~Vries et~al.(2019)de~Vries, van Cranenburgh, Bisazza, Caselli, van
  Noord, and Nissim}]{abs-1912-09582}
Wietse de~Vries, Andreas van Cranenburgh, Arianna Bisazza, Tommaso Caselli,
  Gertjan van Noord, and Malvina Nissim. 2019.
\newblock \href {http://arxiv.org/abs/1912.09582} {Bertje: {A} dutch {BERT}
  model}.
\newblock \emph{CoRR}, abs/1912.09582.

\bibitem[{Devlin et~al.(2018{\natexlab{a}})Devlin, Chang, Lee, and
  Toutanova}]{bert}
Jacob Devlin, Ming{-}Wei Chang, Kenton Lee, and Kristina Toutanova.
  2018{\natexlab{a}}.
\newblock \href {http://arxiv.org/abs/1810.04805} {{{BERT:} Pre-training of
  Deep Bidirectional Transformers for Language Understanding}}.
\newblock \emph{CoRR}, abs/1810.04805.

\bibitem[{Devlin et~al.(2018{\natexlab{b}})Devlin, Chang, Lee, and
  Toutanova}]{mbert}
Jacob Devlin, Ming{-}Wei Chang, Kenton Lee, and Kristina Toutanova.
  2018{\natexlab{b}}.
\newblock \href {http://arxiv.org/abs/1810.04805} {{BERT:} pre-training of deep
  bidirectional transformers for language understanding}.
\newblock \emph{CoRR}, abs/1810.04805.

\bibitem[{Elmadany et~al.(2018)Elmadany, Mubarak, and
  Magdy}]{elmadany2018arsas}
A~Elmadany, Hamdy Mubarak, and Walid Magdy. 2018.
\newblock Arsas: An arabic speech-act and sentiment corpus of tweets.
\newblock \emph{OSACT}, 3:20.

\bibitem[{Feng et~al.(2022)Feng, Yang, Cer, Arivazhagan, and Wang}]{LaBSe}
Fangxiaoyu Feng, Yinfei Yang, Daniel Cer, Naveen Arivazhagan, and Wei Wang.
  2022.
\newblock \href {https://doi.org/10.18653/V1/2022.ACL-LONG.62}
  {Language-agnostic {BERT} sentence embedding}.
\newblock In \emph{Proceedings of the 60th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers), {ACL} 2022, Dublin,
  Ireland, May 22-27, 2022}, pages 878--891. Association for Computational
  Linguistics.

\bibitem[{Inoue et~al.(2021)Inoue, Alhafni, Baimukan, Bouamor, and
  Habash}]{inoue-etal-2021-interplay}
Go~Inoue, Bashar Alhafni, Nurpeiis Baimukan, Houda Bouamor, and Nizar Habash.
  2021.
\newblock The interplay of variant, size, and task type in {A}rabic pre-trained
  language models.
\newblock In \emph{Proceedings of the Sixth Arabic Natural Language Processing
  Workshop}, Kyiv, Ukraine (Online). Association for Computational Linguistics.

\bibitem[{Itamar and Itai(2008)}]{ItamarI08}
Einav Itamar and Alon Itai. 2008.
\newblock \href
  {http://www.lrec-conf.org/proceedings/lrec2008/summaries/76.html} {{Using
  Movie Subtitles for Creating a Large-Scale Bilingual Corpora}}.
\newblock In \emph{Proceedings of the International Conference on Language
  Resources and Evaluation, {LREC} 2008, 26 May - 1 June 2008, Marrakech,
  Morocco}. European Language Resources Association.

\bibitem[{Kiritchenko et~al.(2016)Kiritchenko, Mohammad, and
  Salameh}]{kiritchenko-etal-2016-semeval}
Svetlana Kiritchenko, Saif Mohammad, and Mohammad Salameh. 2016.
\newblock \href {https://doi.org/10.18653/v1/S16-1004} {{S}em{E}val-2016 task
  7: Determining sentiment intensity of {E}nglish and {A}rabic phrases}.
\newblock In \emph{Proceedings of the 10th International Workshop on Semantic
  Evaluation ({S}em{E}val-2016)}, pages 42--51, San Diego, California.
  Association for Computational Linguistics.

\bibitem[{Lan et~al.(2020)Lan, Chen, Goodman, Gimpel, Sharma, and
  Soricut}]{albert}
Zhenzhong Lan, Mingda Chen, Sebastian Goodman, Kevin Gimpel, Piyush Sharma, and
  Radu Soricut. 2020.
\newblock \href {https://openreview.net/forum?id=H1eA7AEtvS} {{ALBERT:} {A}
  lite {BERT} for self-supervised learning of language representations}.
\newblock In \emph{8th International Conference on Learning Representations,
  {ICLR} 2020, Addis Ababa, Ethiopia, April 26-30, 2020}. OpenReview.net.

\bibitem[{Le et~al.(2020)Le, Vial, Frej, Segonne, Coavoux, Lecouteux, Allauzen,
  Crabb{\'{e}}, Besacier, and Schwab}]{LeVFSCLACBS20}
Hang Le, Lo{\"{\i}}c Vial, Jibril Frej, Vincent Segonne, Maximin Coavoux,
  Benjamin Lecouteux, Alexandre Allauzen, Beno{\^{\i}}t Crabb{\'{e}}, Laurent
  Besacier, and Didier Schwab. 2020.
\newblock \href {https://aclanthology.org/2020.lrec-1.302/} {Flaubert:
  Unsupervised language model pre-training for french}.
\newblock In \emph{Proceedings of The 12th Language Resources and Evaluation
  Conference, {LREC} 2020, Marseille, France, May 11-16, 2020}, pages
  2479--2490. European Language Resources Association.

\bibitem[{Liu et~al.(2019)Liu, Ott, Goyal, Du, Joshi, Chen, Levy, Lewis,
  Zettlemoyer, and Stoyanov}]{roberta}
Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer
  Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov. 2019.
\newblock \href {http://arxiv.org/abs/1907.11692} {{RoBERTa: {A} Robustly
  Optimized {BERT} Pretraining Approach}}.
\newblock \emph{CoRR}, abs/1907.11692.

\bibitem[{Malmsten et~al.(2020)Malmsten, B{\"{o}}rjeson, and
  Haffenden}]{abs-2007-01658}
Martin Malmsten, Love B{\"{o}}rjeson, and Chris Haffenden. 2020.
\newblock \href {http://arxiv.org/abs/2007.01658} {Playing with words at the
  national library of sweden - making a swedish {BERT}}.
\newblock \emph{CoRR}, abs/2007.01658.

\bibitem[{Martin et~al.(2020)Martin, Muller, Su{\'{a}}rez, Dupont, Romary,
  de~la Clergerie, Seddah, and Sagot}]{MartinMSDRCSS20}
Louis Martin, Benjamin Muller, Pedro Javier~Ortiz Su{\'{a}}rez, Yoann Dupont,
  Laurent Romary, {\'{E}}ric de~la Clergerie, Djam{\'{e}} Seddah, and
  Beno{\^{\i}}t Sagot. 2020.
\newblock \href {https://doi.org/10.18653/V1/2020.ACL-MAIN.645} {Camembert: a
  tasty french language model}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics, {ACL} 2020, Online, July 5-10, 2020}, pages
  7203--7219. Association for Computational Linguistics.

\bibitem[{Mubarak et~al.(2020)Mubarak, Darwish, Magdy, Elsayed, and
  Al-Khalifa}]{mubarak-etal-2020-overview}
Hamdy Mubarak, Kareem Darwish, Walid Magdy, Tamer Elsayed, and Hend Al-Khalifa.
  2020.
\newblock \href {https://aclanthology.org/2020.osact-1.7} {Overview of {OSACT}4
  {A}rabic offensive language detection shared task}.
\newblock In \emph{Proceedings of the 4th Workshop on Open-Source Arabic
  Corpora and Processing Tools, with a Shared Task on Offensive Language
  Detection}, pages 48--52, Marseille, France. European Language Resource
  Association.

\bibitem[{Mubarak et~al.(2021)Mubarak, Hassan, and
  Abdelali}]{mubarak-etal-2021-adult}
Hamdy Mubarak, Sabit Hassan, and Ahmed Abdelali. 2021.
\newblock \href {https://aclanthology.org/2021.wanlp-1.14} {Adult content
  detection on {A}rabic {T}witter: Analysis and experiments}.
\newblock In \emph{Proceedings of the Sixth Arabic Natural Language Processing
  Workshop}, pages 136--144, Kyiv, Ukraine (Virtual). Association for
  Computational Linguistics.

\bibitem[{Murtadha et~al.(2022)Murtadha, Pan, Wen, Su, Zhang, and
  Liu}]{bert-asc}
Ahmed Murtadha, Shengfeng Pan, Bo~Wen, Jianlin Su, Wenze Zhang, and Yunfeng
  Liu. 2022.
\newblock \href {https://doi.org/10.48550/ARXIV.2203.11702} {{BERT-ASC:}
  auxiliary-sentence construction for implicit aspect learning in sentiment
  analysis}.
\newblock \emph{CoRR}, abs/2203.11702.

\bibitem[{Murtadha et~al.(2024)Murtadha, Wen, Ao, Pan, Su, Cao, and
  Liu}]{AhmedWAPSCL24}
Ahmed Murtadha, Bo~Wen, Luo Ao, Shengfeng Pan, Jianlin Su, Xinxin Cao, and
  Yunfeng Liu. 2024.
\newblock \href {https://doi.org/10.1016/J.INS.2024.120160} {Towards robust
  learning with noisy and pseudo labels for text classification}.
\newblock \emph{Inf. Sci.}, 661:120160.

\bibitem[{Nabil et~al.(2015)Nabil, Aly, and Atiya}]{NabilAA15}
Mahmoud Nabil, Mohamed~A. Aly, and Amir~F. Atiya. 2015.
\newblock \href {https://doi.org/10.18653/V1/D15-1299} {{ASTD:} arabic
  sentiment tweets dataset}.
\newblock In \emph{Proceedings of the 2015 Conference on Empirical Methods in
  Natural Language Processing, {EMNLP} 2015, Lisbon, Portugal, September 17-21,
  2015}, pages 2515--2519. The Association for Computational Linguistics.

\bibitem[{Nguyen and Nguyen(2020)}]{NguyenN20}
Dat~Quoc Nguyen and Anh~Tuan Nguyen. 2020.
\newblock \href {https://doi.org/10.18653/V1/2020.FINDINGS-EMNLP.92} {Phobert:
  Pre-trained language models for vietnamese}.
\newblock In \emph{Findings of the Association for Computational Linguistics:
  {EMNLP} 2020, Online Event, 16-20 November 2020}, volume {EMNLP} 2020 of
  \emph{Findings of {ACL}}, pages 1037--1042. Association for Computational
  Linguistics.

\bibitem[{OpenAI(2023)}]{gpt4}
OpenAI. 2023.
\newblock \href {https://doi.org/10.48550/ARXIV.2303.08774} {{GPT-4} technical
  report}.
\newblock \emph{CoRR}, abs/2303.08774.

\bibitem[{Pyysalo et~al.(2021)Pyysalo, Kanerva, Virtanen, and
  Ginter}]{PyysaloKVG21}
Sampo Pyysalo, Jenna Kanerva, Antti Virtanen, and Filip Ginter. 2021.
\newblock \href {https://aclanthology.org/2021.nodalida-main.1/} {Wikibert
  models: Deep transfer learning for many languages}.
\newblock In \emph{Proceedings of the 23rd Nordic Conference on Computational
  Linguistics, NoDaLiDa 2021, Reykjavik, Iceland (Online), May 31 - June 2,
  2021}, pages 1--10. Link{\"{o}}ping University Electronic Press, Sweden.

\bibitem[{Raffel et~al.(2020)Raffel, Shazeer, Roberts, Lee, Narang, Matena,
  Zhou, Li, and Liu}]{t5}
Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael
  Matena, Yanqi Zhou, Wei Li, and Peter~J. Liu. 2020.
\newblock \href {http://jmlr.org/papers/v21/20-074.html} {Exploring the limits
  of transfer learning with a unified text-to-text transformer}.
\newblock \emph{J. Mach. Learn. Res.}, 21:140:1--140:67.

\bibitem[{Safaya et~al.(2020)Safaya, Abdullatif, and Yuret}]{SafayaAY20}
Ali Safaya, Moutasem Abdullatif, and Deniz Yuret. 2020.
\newblock \href {https://doi.org/10.18653/V1/2020.SEMEVAL-1.271} {{KUISAIL} at
  semeval-2020 task 12: {BERT-CNN} for offensive speech identification in
  social media}.
\newblock In \emph{Proceedings of the Fourteenth Workshop on Semantic
  Evaluation, SemEval@COLING 2020, Barcelona (online), December 12-13, 2020},
  pages 2054--2059. International Committee for Computational Linguistics.

\bibitem[{Su et~al.(2024)Su, Ahmed, Lu, Pan, Bo, and Liu}]{rope}
Jianlin Su, Murtadha Ahmed, Yu~Lu, Shengfeng Pan, Wen Bo, and Yunfeng Liu.
  2024.
\newblock \href {https://doi.org/10.1016/J.NEUCOM.2023.127063} {Roformer:
  Enhanced transformer with rotary position embedding}.
\newblock \emph{Neurocomputing}, 568:127063.

\bibitem[{Talafha et~al.(2020)Talafha, Ali, Za'ter, Seelawi, Tuffaha, Samir,
  Farhan, and Al{-}Natsheh}]{mdbert}
Bashar Talafha, Mohammad Ali, Muhy~Eddin Za'ter, Haitham Seelawi, Ibraheem
  Tuffaha, Mostafa Samir, Wael Farhan, and Hussein~T. Al{-}Natsheh. 2020.
\newblock \href {https://www.aclweb.org/anthology/2020.wanlp-1.10/}
  {Multi-dialect arabic {BERT} for country-level dialect identification}.
\newblock In \emph{Proceedings of the Fifth Arabic Natural Language Processing
  Workshop, WANLP@COLING 2020, Barcelona, Spain (Online), December 12, 2020},
  pages 111--118. Association for Computational Linguistics.

\bibitem[{Touvron et~al.(2023)Touvron, Martin, Stone, Albert, Almahairi,
  Babaei, Bashlykov, Batra, Bhargava, Bhosale, Bikel, Blecher, Canton{-}Ferrer,
  Chen, Cucurull, Esiobu, Fernandes, Fu, Fu, Fuller, Gao, Goswami, Goyal,
  Hartshorn, Hosseini, Hou, Inan, Kardas, Kerkez, Khabsa, Kloumann, Korenev,
  Koura, Lachaux, Lavril, Lee, Liskovich, Lu, Mao, Martinet, Mihaylov, Mishra,
  Molybog, Nie, Poulton, Reizenstein, Rungta, Saladi, Schelten, Silva, Smith,
  Subramanian, Tan, Tang, Taylor, Williams, Kuan, Xu, Yan, Zarov, Zhang, Fan,
  Kambadur, Narang, Rodriguez, Stojnic, Edunov, and Scialom}]{llam2}
Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine
  Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale,
  Dan Bikel, Lukas Blecher, Cristian Canton{-}Ferrer, Moya Chen, Guillem
  Cucurull, David Esiobu, Jude Fernandes, Jeremy Fu, Wenyin Fu, Brian Fuller,
  Cynthia Gao, Vedanuj Goswami, Naman Goyal, Anthony Hartshorn, Saghar
  Hosseini, Rui Hou, Hakan Inan, Marcin Kardas, Viktor Kerkez, Madian Khabsa,
  Isabel Kloumann, Artem Korenev, Punit~Singh Koura, Marie{-}Anne Lachaux,
  Thibaut Lavril, Jenya Lee, Diana Liskovich, Yinghai Lu, Yuning Mao, Xavier
  Martinet, Todor Mihaylov, Pushkar Mishra, Igor Molybog, Yixin Nie, Andrew
  Poulton, Jeremy Reizenstein, Rashi Rungta, Kalyan Saladi, Alan Schelten, Ruan
  Silva, Eric~Michael Smith, Ranjan Subramanian, Xiaoqing~Ellen Tan, Binh Tang,
  Ross Taylor, Adina Williams, Jian~Xiang Kuan, Puxin Xu, Zheng Yan, Iliyan
  Zarov, Yuchen Zhang, Angela Fan, Melanie Kambadur, Sharan Narang,
  Aur{\'{e}}lien Rodriguez, Robert Stojnic, Sergey Edunov, and Thomas Scialom.
  2023.
\newblock \href {https://doi.org/10.48550/ARXIV.2307.09288} {Llama 2: Open
  foundation and fine-tuned chat models}.
\newblock \emph{CoRR}, abs/2307.09288.

\bibitem[{Virtanen et~al.(2019)Virtanen, Kanerva, Ilo, Luoma, Luotolahti,
  Salakoski, Ginter, and Pyysalo}]{abs-1912-07076}
Antti Virtanen, Jenna Kanerva, Rami Ilo, Jouni Luoma, Juhani Luotolahti, Tapio
  Salakoski, Filip Ginter, and Sampo Pyysalo. 2019.
\newblock \href {http://arxiv.org/abs/1912.07076} {Multilingual is not enough:
  {BERT} for finnish}.
\newblock \emph{CoRR}, abs/1912.07076.

\bibitem[{Xue et~al.(2021)Xue, Constant, Roberts, Kale, Al{-}Rfou, Siddhant,
  Barua, and Raffel}]{mT5}
Linting Xue, Noah Constant, Adam Roberts, Mihir Kale, Rami Al{-}Rfou, Aditya
  Siddhant, Aditya Barua, and Colin Raffel. 2021.
\newblock \href {https://doi.org/10.18653/V1/2021.NAACL-MAIN.41} {mt5: {A}
  massively multilingual pre-trained text-to-text transformer}.
\newblock In \emph{Proceedings of the 2021 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, {NAACL-HLT} 2021, Online, June 6-11, 2021}, pages 483--498.
  Association for Computational Linguistics.

\bibitem[{Zhu et~al.(2015)Zhu, Kiros, Zemel, Salakhutdinov, Urtasun, Torralba,
  and Fidler}]{ZhuKZSUTF15}
Yukun Zhu, Ryan Kiros, Richard~S. Zemel, Ruslan Salakhutdinov, Raquel Urtasun,
  Antonio Torralba, and Sanja Fidler. 2015.
\newblock \href {https://doi.org/10.1109/ICCV.2015.11} {Aligning books and
  movies: Towards story-like visual explanations by watching movies and reading
  books}.
\newblock In \emph{2015 {IEEE} International Conference on Computer Vision,
  {ICCV} 2015, Santiago, Chile, December 7-13, 2015}, pages 19--27. {IEEE}
  Computer Society.

\end{thebibliography}

\pdfoutput=1
\documentclass[11pt]{article}

\usepackage[]{acl}
\usepackage{times}
\usepackage{adjustbox}
\usepackage{multirow}
\usepackage{latexsym}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{makecell}
\usepackage{amssymb}

\usepackage{mathabx}
\usepackage[T1]{fontenc}
\usepackage[utf8]{inputenc}

\usepackage{listings}
\usepackage{xcolor}

\usepackage{microtype}

\title{\textbf{AlcLaM}: Arabic Dialectal  Language Model}
\author{
		\adjustbox{width=\linewidth}{
	Murtadha Ahmed\textsuperscript{*}, Saghir Alfasly\textsuperscript{\dag},  Bo Wen\textsuperscript{*},  Jamaal Qasem\textsuperscript{\ddag}, Mohammed Ahmed\textsuperscript{\S}, Yunfeng Liu\textsuperscript{*}}\\
		\textsuperscript{*}Zhuiyi AI Lab, China, 
		\textsuperscript{\dag}Mayo Clinic, USA, \textsuperscript{\ddag}Dongbei University of Finance and  \\
		Economics, China,
		\textsuperscript{\S}Northwestern Polytechnical University, China \\
	\adjustbox{width=\linewidth}{{\textsuperscript{*}\tt \{a.murtadha,brucewen,glenliu\}@wezhuiyi.com}, {\textsuperscript{\dag}\tt alfasly.saghir@mayo.edu}}\\
	{\textsuperscript{\ddag}\tt ja.qasem@dufe.edu.cn},
	{\textsuperscript{\S}\tt majeedi@mail.nwpu.edu.cn}
}

\begin{document}
\maketitle
%\begin{abstract}
%-trained language models (LMs) are fundamental components of many contemporary natural language processing (NLP) systems. While multilingual models address a broad spectrum of languages, they face challenges such as high inference costs and limited diversity in non-English training data. Arabic LMs predominantly rely on Modern Standard Arabic, significantly reducing their efficacy with Arabic dialects. To address this, we first build an Arabic dialectal corpus of 3,372,744 sentences sourced from social media. Then, we expand the vocabulary and retrained a BERT model from scratch using this extensive corpus. Despite our AlcLaM is only trained in 13 GB text, due to computational resources limitation,  which is 7.8\%, 10.2\% and 21.3\$ of the training data used by  CAMeLBERT, MDBERT and ArBERT, respectively, AlcLaM can surprisingly achieve better performance across a range of Arabic NLP task. The model will be made publicly available to foster further research and application improvements.
%\end{abstract}

\begin{abstract}
	Pre-trained Language Models (PLMs) are integral to many modern natural language processing (NLP) systems. Although multilingual models cover a wide range of languages, they often grapple with challenges like high inference costs and a lack of diverse non-English training data. Arabic-specific PLMs are trained predominantly on modern standard Arabic, which compromises their performance on regional dialects. To tackle this, we construct an Arabic dialectal corpus comprising 3.4M sentences gathered from social media platforms. We utilize this corpus to expand the vocabulary and retrain a BERT-based model from scratch. Named AlcLaM, our model was trained using only 13 GB of text, which represents a fraction of the data used by existing models such as CAMeL, MARBERT, and ArBERT,  compared to 7.8\%, 10.2\%, and 21.3\%, respectively. Remarkably, AlcLaM demonstrates superior performance on a variety of Arabic NLP tasks despite the limited training data. AlcLaM is available at:
	\href{https://github.com/amurtadha/Alclam}{https://github.com/amurtadha/Alclam}.
%	 to encourage further academic research and enhancements in practical applications.
\end{abstract}

\section{Introduction}
Pre-trained Language Models (PLMs) utilizing self-supervised learning techniques, such as BERT \cite{bert} and RoBERTa \cite{roberta}, have become pivotal in advancing the field of natural language processing (NLP) through transfer learning. These models have significantly enhanced performance across a variety of NLP tasks by leveraging vast amounts of textual data and extensive computational resources. However, the necessity for large corpora and the substantial computational demand, often requiring weeks of training time \cite{conneau-etal-2020-unsupervised,t5,abs-2001-09977}, has primarily confined the development of such models to the English language and a few other major languages.

This limitation has sparked an increased interest in creating multilingual models capable of understanding and processing multiple languages simultaneously. Innovations such as mBERT \cite{bert}, XLM-RoBERTa \cite{conneau-etal-2020-unsupervised} and  LaBSE \cite{LaBSe} aim to address this gap. Despite these efforts, the performance of these multilingual models typically lags behind their monolingual counterparts. This discrepancy is largely due to smaller, language-specific vocabularies and less comprehensive language-specific datasets \cite{abs-1912-07076,antoun2020arabert,DadasPP20,abs-1912-09582,abs-2007-01658,NguyenN20}.

Furthermore, while languages with similar structures and vocabularies may benefit from shared representations \cite{conneau-etal-2020-unsupervised}, this advantage does not extend to languages such as Arabic. Arabic's unique morphological and syntactic structures share little in common with the morphosyntactic frameworks of more abundantly represented Latin-based languages. To address this, various Arabic-specific PLMs have been developed, including AraBERT \cite{antoun2020arabert}, ArBERT \cite{Abdul-MageedEN20}, and CAMeL \cite{inoue-etal-2021-interplay}. These models significantly enhance Arabic NLP tasks over multilingual models. However, they are predominantly trained on Modern Standard Arabic (MSA) datasets. This focus on MSA introduces two primary limitations: first, there is reduced recognition of dialectal tokens, which vary widely across different Arabic-speaking regions; second, there is a biased weighting towards MSA tokens in the models, which may not accurately reflect the linguistic nuances present in everyday Arabic usage.

In this paper, we first introduce a new corpus of 3,372,744 Arabic dialectal texts, meticulously sourced from social media platforms such as YouTube and Facebook. Second, we outline the procedure for pretraining the  transformer model \cite{bert} specifically for the Arabic language, which we dub AlcLaM. Note that we only train AlcLaM on 13GB text due to computational resources limitation.   Finally, we assess AlcLaM's performance on three Arabic NLU downstream tasks, each distinct in nature: (i) Arabic Dialect Identification (DID), (ii) Sentiment Analysis (SA), and (iii) Hate Speech and Offensive Language Detection.   Despite the limited training data, our experimental results demonstrate that AlcLaM attains state-of-the-art results on most datasets, surpassing several baseline models, including previous multilingual and single-language approaches.

%In brief, our contributions are two-fold. (1) Construction of a large corpus of Arabic dialects sourced from social media. (2) Development of AlcLaM, an Arabic PLM tailored to accommodate the diversity and complexity of Arabic dialect text.

\iftrue
In summary, our contributions are twofold.
\begin{itemize}
\item We constructed a massive corpus of Arabic dialects, derived from the content and comments on Arabic pages on Facebook and videos from Arabic-speaking YouTubers. This corpus represents a rich variety of regional dialects and everyday language usage that has been underrepresented in previous models.
\item We developed an Arabic pre-trained language model, namely AlcLaM, specifically optimized to handle the diversity and complexity of Arabic dialects based on the newly created corpus, enhancing its applicability across a wider range of NLP tasks involving Arabic text.

%\item Constructed a vast Arabic dialect corpus from Facebook pages and Arabic-speaking YouTubers, capturing diverse regional dialects and everyday language usage.
%\item Developed AlcLaM, an Arabic pre-trained language model tailored to handle Arabic dialect diversity, expanding its utility across diverse NLP tasks.
\end{itemize}
\fi

\section{Related Work}
Pre-trained language models (PLMs) using a self-supervised masking objective, such as BERT \cite{bert} and RoBERTa \cite{roberta}, have significantly advanced NLP. These models have multilingual versions, including  mBERT \cite{bert}, XLM-RoBERTa \cite{conneau-etal-2020-unsupervised} and  LaBSE \cite{LaBSe}. Additionally, models featuring different objectives or architectures, such as ALBERT \cite{albert}, T5 \cite{t5}, its multilingual variant mT5 \cite{mT5}, and GPT-3 \cite{gpt3}, LLaMA \cite{llam2},  PaLM \cite{palm},  GPT-4 \cite{gpt4}, and  RoFormer \cite{rope}  have been introduced. 

Non-English PLMs have also been developed. These include Bertje for Dutch \cite{abs-1912-09582}, CamemBERT \cite{MartinMSDRCSS20} and FlauBERT \cite{LeVFSCLACBS20} for French, PhoBERT for Vietnamese \cite{NguyenN20}, as well as models for Finnish by \citet{abs-1912-07076}, for Polish by \citet{DadasPP20}, and for Swedish by \citet{abs-2007-01658}. \citet{PyysaloKVG21} have created monolingual LMs using Wikipedia data for 42 languages. For Arabic, MSA-based PLMs includes AraBERT \cite{antoun2020arabert} ArabicBERT \cite{SafayaAY20}, ArBERT \cite{Abdul-MageedEN20}. 
Another line of research involves pre-training models on a combination of MSA and dialectal data, such as MDBERT \cite{Abdul-MageedEN20} and CAMeL \cite{mubarak-etal-2021-adult}. Our contributions to this field include a comprehensive Arabic dialectal corpus spanning various dialects and the development of an Arabic PLM. Our model, named AlcLaM, enhances the representation of linguistic diversity in Arabic NLP.

\section{Methodology}
In this paper, we develop AlcLaM, an Arabic dialect language representation model that enhances the performance on several Arabic NLP tasks. This model builds upon the BERT architecture, a stacked Bidirectional Transformer Encoder \cite{bert}. Recognized as the foundation for many state-of-the-art results in various NLP tasks across multiple languages, BERT's architecture has proven highly effective.  
%For AlcLaM, we adopt the RoBERTa-base configuration \cite{roberta}, which includes 12 encoder blocks, 768 hidden dimensions, 12 attention heads, a maximum sequence length of 512, and approximately 125 million parameters. 
Below, we detail  the dialectal corpus  used for AlcLaM's pretraining, the pretraining setup, and the fine-tuning process.

\subsection{Arabic Dialectal Corpus.}
The original BERT model was trained on a corpus comprising 3.3 billion words extracted from English Wikipedia and the Book Corpus \cite{ZhuKZSUTF15}.
%The original RoBERTa model was trained on a vast dataset consisting of 160GB of uncompressed text sourced from various corpora, including CC-NEWS, the Book Corpus (Zhu et al., 2015), OPENWEBTEXT (Gokaslan and Cohen, 2019), and STORIES (Trinh and Le, 2018). 
Due to the comparatively smaller size of Arabic Wikipedia dumps in comparison to English ones, we opted to utilize Arabic text from the English-Arabic bilingual corpora of opensubtitles \footnote{\href{https://opus.nlpl.eu/OpenSubtitles/}{opensubtitles}} \cite{ItamarI08}. 

It is noteworthy that publicly available Arabic corpora are heavily dominated by MSA, while social media and online reviews predominantly feature Arabic dialects. This creates a bias towards MSA tokens in Arabic PLMs, potentially leading to tokenizers failing to recognize a significant portion of dialectal vocabulary. To address this, we manually scraped Arabic texts from social media platforms. Initially, we scrape posts and comments from popular Arabic YouTube channels and Facebook Pages. However, we observed that many of these comments consisted of verses from the Holy Quran and Hadith, typically written in MSA. Since our focus was on dialectal texts, we trained a binary classifier (MSA-Dialect) to filter out MSA texts. Specifically, we treated all dialectal instances of the MADAR corpus as one class, labeled "Dialect" \cite{BouamorHH19,bert-asc}, and utilized it to fine-tune the CAMeL model \cite{inoue-etal-2021-interplay}, which achieved a remarkable 98\% accuracy. Our final corpus comprises 3,372,744 dialectal sentences with 54,557,408 tokens. To the best of our knowledge, this marks the first attempt to assemble such a comprehensive Arabic dialectal corpus.

%It's important to acknowledge that publicly accessible Arabic corpora predominantly feature Modern Standard Arabic (), while Arabic dialects are prevalent in social media and reviews. This creates a bias towards MSA tokens in Arabic pre-trained language models (PLMs), potentially causing tokenizers to overlook a significant portion of dialectal vocabulary. To address this disparity, we manually curated Arabic texts from social media platforms. Initially, we leveraged statistical insights from popular Arabic YouTube channels and Facebook Pages to gather their comments. However, it became apparent that many of these comments consisted of verses from the Holy Quran or possibly hadiths, typically composed in MSA. Given our focus on dialectal texts, we devised a binary classifier (MSA-Dialect) to sift out MSA texts.
%
%Specifically, we treated all dialectal instances of the MADAR corpus as a single class, labeled "Dialect" \cite{bibid}. We then employed this classification to fine-tune the CAMeL model \cite{bibid}, achieving an impressive 98% accuracy. Our final corpus comprises ... sentences and ... tokens. To our knowledge, this represents the inaugural effort to assemble such a comprehensive Arabic dialectal corpus.

\subsection{Model Training}

For AlcLaM, we adhere to the original BERT \cite{bert}. Each training input sequence is generated using whole word masking, where 15\% of the N input tokens are chosen for replacement. These selected tokens undergo replacement as follows: 80\% are substituted with the [MASK] token, 10\% with a random token, and 10\% remain unchanged. Following \citet{roberta}, we exclude the next sentence prediction (NSP) loss from our training process. This decision is based on the observation that removing the NSP loss either matches or slightly improves downstream task performance. We employ the same network configuration as BERT-base: consisting of 12 layers, 768 hidden units, and 12 attention heads, resulting in approximately 125 million parameters.
During training, we utilize a batch size of 64 sequences and set a maximum sequence length of 128 tokens and 5 training epochs. 
%This training duration roughly equates to 42 epochs over the corpus of 6.5 billion tokens.
Throughout training, we set the learning rate to $5e-5$.

\subsection{Fine-tuning}
To fine-tune AlcLaM for sequence classification, we utilize the final hidden state of the first token, corresponding to the embedding of the special “[CLS]” token that is prepended to the beginning of each sentence \cite{AhmedWAPSCL24}. A simple feed-forward layer with a Softmax activation function is added to compute the probability distribution over the predicted output classes. During fine-tuning, both the classifier and the pre-trained model weights are jointly trained to maximize the log-probability of the correct class \cite{rnt}.

\section{Empirical Evaluation}
\subsection{Datasets}
We evaluated AlcLaM on the following datasets that cover various NLP tasks in Arabic. Sentiment analysis (SemEval 2017 task 4 \cite{kiritchenko-etal-2016-semeval}, ASAD \cite{abs-2011-00578}, ASTD \cite{NabilAA15}, ArSAS \cite{elmadany2018arsas}, LABR \cite{AlyA13}), offensive language detection (Adult \cite{mubarak-etal-2021-adult}, Offensive and HateSpeech \cite{mubarak-etal-2020-overview}), dialect identification (MADAR-6, MADAR-26\cite{BouamorHH19} and NADI \cite{abdul-mageed-etal-2020-nadi}). For experiments, MADAR-2 and MADAR-9 are derived from MADAR-26. MADAR-2 is binary (MSA-dialect), while MADAR-9 categorizes dialects into 9 regions: Yemen, MSA, Maghreb, Nile Egypt, Libya, Gulf, Nile Sudan, Iraq, and Levant.

\subsection{Baselines}
%We compare our AlcLaM model with various multilingual PLMs, such as LaBSE \cite{LaBSe} and mBERT \cite{mbert}, as well as Arabic-specific PLMs like MDBERT \cite{mdbert}, AraBERT \cite{antoun2020arabert}, ArBERT, and MARBERT \cite{Abdul-MageedEN20}. Additionally, we include CAMeL \cite{inoue-etal-2021-interplay} in our comparison. It is worth noting that, to the best of our knowledge, only CAMeL and MARBERT have been trained on data encompassing both MSA and diverse Arabic dialects.

%We compare our AlcLaM to (I) multilingual PLMs, such as LaBSE \cite{LaBSe} and mBERT \cite{mbert}; (II) MSA-based Arabic PLMs, AraBERT \cite{antoun2020arabert}, ArBERT  \cite{Abdul-MageedEN20}, and  MDBERT \cite{mdbert},  as well as (III) MSA-Dialect-based PLMs like,  MARBERT \cite{Abdul-MageedEN20} and   CAMeL \cite{inoue-etal-2021-interplay}. It is worth noting that, to the best of our knowledge, only CAMeL and MARBERT have been trained on data encompassing both MSA and diverse Arabic dialects.

We compare our AlcLaM model with:
\begin{enumerate}
\item  Multilingual PLMs like mBERT \cite{mbert} and  LaBSE \cite{LaBSe};

\item MSA-based Arabic PLMs such as AraBERT \cite{antoun2020arabert} and ArBERT \cite{Abdul-MageedEN20};

\item MSA-Dialect-based PLMs, including  MdBERT \cite{mdbert}, and MARBERT \cite{Abdul-MageedEN20} and CAMeL \cite{inoue-etal-2021-interplay}.

\end{enumerate}

%It is noteworthy that, to our knowledge, only MdBERT, CAMeL and MARBERT have been trained on data encompassing both MSA and a diverse range of Arabic dialects.

\subsection{Results}
%Note that for each datasets, we report the averaged results of five runs with different randomization to significant test. We report the results of various Arabic NLP tasks in Table \ref{tab:acc} and Table \ref{tab:f1} in terms of the accuracy and F1 metrics, respectively, from which we have made the following observations. (1) Multilingual models (e.g., mBERT, LaBSE) are outperformed by Arabic models pre-trained with larger vocabulary and bigger language-specific datasets and this align with the findings of \citet{Abdul-MageedEN20}. (2) The models that involve dialectal data (e.g., MARBERT, CAMeL) during the pre-training surprisingly achieve better performance not only  Arabic Dialect Identification but across a range of Arabic NLP Task. This finding may kindly argue with the claim that similarity in Arabic dialects may reflect positively on other tasks rather than ADI. The experimental results demonstrates the need to leverage more dialectal information during the training. We believe that this performance is reasonable as the tokenizers may recognize more dialectal tokens, which are consider as unknown in the other models. (3) Despite that our model is trained on less MSA text and with less steps, due to computational resource limitation, it outperforms it alternatives in most tasks and achieve competitive performance on the other. It is true that the improvements on other tasks  is   modest compare to its improvements on ADI, yet given the recognized challenge of Arabic language, these improvements can deemed significant. 

\begin{table*}[t]
	\centering
	\adjustbox{width=\linewidth}{		
		\begin{tabular}{llllllllllll}  
			\hline

			&\multirow{2}{*}{Dataset}&\multicolumn{2}{c}{Multilingual PLMs}&&\multicolumn{2}{c}{MSA-based PLMs}&&\multicolumn{4}{c}{MSA-Dialect-based PLMs}\\\cline{3-4}\cline{6-7}\cline{9-12}
			&&mBERT&LaBSE&&AraBERT&ArBERT&&MdBERT&CAMeL&MARBERT&AlcLaM\\
			\hline
			%			\hline \hline  \multicolumn{11}{c}{\textbf{Dialect Identification}}\\ \hline \hline 
			\multirow{5}{*}{\makecell{DID}}
			&	MADAR-2& 72.9 $\pm$ 16.9& 86.6 $\pm$ 0.5&& 87.1 $\pm$ 0.2& 87.1 $\pm$ 0.2&& 86.0 $\pm$ 0.6& 87.5 $\pm$ 1.0& 85.3 $\pm$ 3.8& \textbf{98.2} $\pm$ \textbf{0.1}\\
			&	MADAR-6& 91.3 $\pm$ 0.1& 91.1 $\pm$ 0.2&& 91.6 $\pm$ 0.1& 91.6 $\pm$ 0.2&& 91.6 $\pm$ 0.0& 92.0 $\pm$ 0.1& 92.2 $\pm$ 0.2& \textbf{93.2} $\pm$ \textbf{0.1}$^*$\\
			&MADAR-9& 75.5 $\pm$ 0.5& 75.7 $\pm$ 0.2&& 76.8 $\pm$ 0.3& 74.5 $\pm$ 4.3&& 75.9 $\pm$ 0.5& 77.5 $\pm$ 0.4& 78.2 $\pm$ 0.3& \textbf{81.9} $\pm$ \textbf{0.3}\\
			&MADAR-26& 60.5 $\pm$ 0.2& 62.0 $\pm$ 0.2&& 62.0 $\pm$ 0.1& 61.7 $\pm$ 0.1&& 60.2 $\pm$ 0.4& 62.9 $\pm$ 0.1& 61.5 $\pm$ 0.4& \textbf{66.3} $\pm$ \textbf{0.1}$^*$\\
			&NADI& 17.6 $\pm$ 0.5& 17.6 $\pm$ 0.5&& 22.6 $\pm$ 0.5& 22.6 $\pm$ 0.5&& 24.9 $\pm$ 0.6& 25.9 $\pm$ 0.5& \textbf{28.6} $\pm$ \textbf{0.8}$^*$& 25.6 $\pm$ 0.6\\
			
			\hline
			\multirow{6}{*}{\makecell{SA}}
			%			\hline \hline  \multicolumn{11}{c}{\textbf{Sentiment Analysis}}\\ \hline \hline 
			&SemEval& 51.3 $\pm$ 1.3& 64.2 $\pm$ 0.7&& 65.4 $\pm$ 0.5& 64.4 $\pm$ 0.9&& 65.6 $\pm$ 0.3& 67.1 $\pm$ 0.7& 66.4 $\pm$ 0.3& \textbf{69.2} $\pm$ \textbf{0.4}$^*$\\
			&ASAD& 59.8 $\pm$ 0.0& 62.4 $\pm$ 0.0&& 41.3 $\pm$ 0.0& 66.9 $\pm$ 0.0&& \textbf{67.5} $\pm$ \textbf{0.0}& 65.8 $\pm$ 0.0& 66.8 $\pm$ 0.0& 66.7 $\pm$ 0.0\\
			&AJGT& 86.4 $\pm$ 0.3& 92.4 $\pm$ 0.7&& 92.7 $\pm$ 0.3& 92.6 $\pm$ 0.4&& 93.6 $\pm$ 0.0& 93.6 $\pm$ 0.3& 93.7 $\pm$ 0.1& \textbf{95.0} $\pm$ \textbf{0.3}$^*$\\
			&ASTD& 46.3 $\pm$ 1.4& 55.7 $\pm$ 0.4&& 57.5 $\pm$ 2.3& 59.7 $\pm$ 0.1&& 61.9 $\pm$ 0.4& 60.2 $\pm$ 0.2& 61.0 $\pm$ 0.5& \textbf{64.6} $\pm$ \textbf{0.1}$^*$\\
			&LABR& 81.1 $\pm$ 0.0& 85.4 $\pm$ 0.0&& 85.9 $\pm$ 0.0& 85.9 $\pm$ 0.0&& 84.7 $\pm$ 0.0& \textbf{86.3} $\pm$ \textbf{0.0}& 85.0 $\pm$ 0.0& 84.9 $\pm$ 0.0\\
			&ARSAS& 73.2 $\pm$ 0.7& 76.2 $\pm$ 0.6&& 76.8 $\pm$ 0.3& 76.1 $\pm$ 0.2&& 76.3 $\pm$ 0.2& 77.1 $\pm$ 0.3& 76.2 $\pm$ 0.2& \textbf{77.9} $\pm$ \textbf{0.3}$^*$\\
			\hline
			\multirow{3}{*}{\makecell{HSOD}}
			%			\hline \hline  \multicolumn{11}{c}{\textbf{Hate Speech and Offensive Language Detection}}\\ \hline \hline 
			&HateSpeech& 67.9 $\pm$ 1.4& 73.7 $\pm$ 1.1&& 76.4 $\pm$ 1.2& 76.8 $\pm$ 1.4&& 80.0 $\pm$ 0.1& 78.8 $\pm$ 0.6& 80.0 $\pm$ 0.8& \textbf{81.4} $\pm$ \textbf{0.5}$^*$\\
			&Offense& 85.3 $\pm$ 0.5& 87.2 $\pm$ 0.5&& 90.5 $\pm$ 0.4& 90.5 $\pm$ 0.4&& 90.8 $\pm$ 0.2& 89.2 $\pm$ 0.5& 90.8 $\pm$ 0.3& \textbf{91.3} $\pm$ \textbf{0.3}$^*$\\
			&Adult& 87.9 $\pm$ 0.1& 87.2 $\pm$ 0.3&& 88.6 $\pm$ 0.1& 88.4 $\pm$ 0.6&& 88.1 $\pm$ 0.0& 88.6 $\pm$ 0.3& 88.3 $\pm$ 0.1& \textbf{89.3} $\pm$ \textbf{0.3}$^*$\\

			\hline

		\end{tabular}
	}		
	\caption{F1 Score Evaluation of Various Arabic NLP Models. Best scores are highlighted in bold. An asterisk (*) denotes statistical significance, determined by a t-test with a p-value (\( < 0.05 \)). Our AlcLaM not only excels in DID task but also shows improvements in most other tasks. This performance is expected as most Arabic NLP datasets are collected from social media, which is dominated by dialectal expressions.}
	\label{tab:f1}	
\end{table*}

\begin{table*}[!htbp]
	\centering
	\adjustbox{width=\linewidth}{		
		\begin{tabular}{llllllllllll}  
			\hline

			&\multirow{2}{*}{Dataset}&\multicolumn{2}{c}{Multilingual PLMs}&&\multicolumn{2}{c}{MSA-based PLMs}&&\multicolumn{4}{c}{MSA-Dialect-based PLMs}\\\cline{3-4}\cline{6-7}\cline{9-12}
			&&mBERT&LaBSE&&AraBERT&ArBERT&&MdBERT&CAMeL&MARBERT&AlcLaM\\
			%			&mBERT&LaBSE&AraBERT&ArBERT&MdBERT&CAMeL&MARBERT&AlcLaM\\
			%			\hline
			%			\item  Multilingual PLMs like LaBSE \cite{LaBSe} and mBERT \cite{mbert};
			%			
			%			\item MSA-based Arabic PLMs such as AraBERT \cite{antoun2020arabert}, ArBERT \cite{Abdul-MageedEN20}, and MDBERT \cite{mdbert};
			%			
			%			\item MSA-Dialect-based PLMs such as MARBERT \cite{Abdul-MageedEN20} and CAMeL \cite{inoue-etal-2021-interplay}.
			%			
			\hline
			\multirow{3}{*}{\makecell{DID}}
			%			\hline \hline  \multicolumn{11}{c}{\textbf{Dialect Identification}}\\ \hline \hline 
			&MADAR-2& 97.3 $\pm$ 0.8& 98.0 $\pm$ 0.1&& 98.1 $\pm$ 0.0& 98.1 $\pm$ 0.0&& 98.0 $\pm$ 0.1& 98.1 $\pm$ 0.1& 97.2 $\pm$ 0.7& \textbf{99.7} $\pm$ \textbf{0.0}\\
			&MADAR-6& 91.3 $\pm$ 0.1& 91.1 $\pm$ 0.2&& 91.6 $\pm$ 0.1& 91.6 $\pm$ 0.2&& 91.6 $\pm$ 0.0& 92.0 $\pm$ 0.1& 92.2 $\pm$ 0.2& \textbf{93.2} $\pm$ \textbf{0.1}$^*$\\
			&MADAR-9& 78.5 $\pm$ 0.5& 79.1 $\pm$ 0.1&& 80.4 $\pm$ 0.2& 77.7 $\pm$ 3.6&& 79.1 $\pm$ 0.5& 80.5 $\pm$ 0.2& 81.1 $\pm$ 0.3& \textbf{83.4} $\pm$ \textbf{0.4}\\
			&MADAR-26& 60.6 $\pm$ 0.2& 61.9 $\pm$ 0.2&& 61.9 $\pm$ 0.1& 61.7 $\pm$ 0.2&& 60.1 $\pm$ 0.3& 62.9 $\pm$ 0.2& 61.3 $\pm$ 0.3& \textbf{66.1} $\pm$ \textbf{0.2}$^*$\\
			&NADI& 33.4 $\pm$ 0.6& 33.4 $\pm$ 0.6&& 38.9 $\pm$ 1.7& 38.9 $\pm$ 1.7&& 41.9 $\pm$ 1.9& 42.7 $\pm$ 1.6& \textbf{47.3} $\pm$ \textbf{0.1}$^*$& 46.6 $\pm$ 1.0\\
			\hline
			\multirow{3}{*}{\makecell{SA}}
			%			\hline \hline  \multicolumn{11}{c}{\textbf{Sentiment Analysis}}\\ \hline \hline 
			&SemEval& 53.4 $\pm$ 1.5& 65.0 $\pm$ 0.6&& 66.1 $\pm$ 0.5& 65.1 $\pm$ 0.8&& 66.1 $\pm$ 0.3& 68.0 $\pm$ 0.3& 66.9 $\pm$ 0.3& \textbf{69.5} $\pm$ \textbf{0.3}$^*$\\
			&ASAD& 74.6 $\pm$ 0.0& 75.2 $\pm$ 0.0&& 70.6 $\pm$ 0.0& 78.4 $\pm$ 0.0&& 77.6 $\pm$ 0.0& 77.0 $\pm$ 0.0& 77.6 $\pm$ 0.0& \textbf{79.5} $\pm$ \textbf{0.0}\\
			&AJGT& 86.4 $\pm$ 0.3& 92.4 $\pm$ 0.7&& 92.8 $\pm$ 0.3& 92.6 $\pm$ 0.4&& 93.6 $\pm$ 0.0& 93.6 $\pm$ 0.3& 93.8 $\pm$ 0.1& \textbf{95.0} $\pm$ \textbf{0.3}$^*$\\
			&ASTD& 46.7 $\pm$ 1.7& 55.6 $\pm$ 0.6&& 57.7 $\pm$ 2.4& 59.7 $\pm$ 0.3&& 62.0 $\pm$ 0.3& 60.1 $\pm$ 0.2& 61.0 $\pm$ 0.3& \textbf{64.9} $\pm$ \textbf{0.1}$^*$\\
			&LABR& 90.4 $\pm$ 0.0& 92.3 $\pm$ 0.0&& 92.8 $\pm$ 0.0& 92.8 $\pm$ 0.0&& 91.9 $\pm$ 0.0& \textbf{93.0} $\pm$ \textbf{0.0}& 92.6 $\pm$ 0.0& 92.6 $\pm$ 0.0\\
			&ARSAS& 74.5 $\pm$ 0.8& 77.2 $\pm$ 0.7&& 77.6 $\pm$ 0.3& 77.0 $\pm$ 0.3&& 77.5 $\pm$ 0.3& 78.0 $\pm$ 0.3& 77.4 $\pm$ 0.4& \textbf{78.6} $\pm$ \textbf{0.5}$^*$\\
			\hline
			\multirow{3}{*}{\makecell{HSOD}}
			%			\hline \hline  \multicolumn{11}{c}{\textbf{Hate Speech and Offensive Language Detection}}\\ \hline \hline 
			&HateSpeech& 75.2 $\pm$ 2.2& 80.0 $\pm$ 0.7&& 80.5 $\pm$ 1.4& 80.8 $\pm$ 1.9&& 84.3 $\pm$ 0.3& 83.3 $\pm$ 0.6& 84.4 $\pm$ 0.4& \textbf{84.6} $\pm$ \textbf{0.7}$^*$\\
			&Offense& 91.7 $\pm$ 0.1& 92.8 $\pm$ 0.4&& 94.5 $\pm$ 0.2& 94.6 $\pm$ 0.4&& 94.6 $\pm$ 0.2& 93.6 $\pm$ 0.2& 94.8 $\pm$ 0.0& \textbf{94.9} $\pm$ \textbf{0.1}$^*$\\
			&Adult& 95.0 $\pm$ 0.0& 94.4 $\pm$ 0.2&& 95.2 $\pm$ 0.1& 94.9 $\pm$ 0.4&& 95.1 $\pm$ 0.1& 95.2 $\pm$ 0.2& 95.1 $\pm$ 0.0& \textbf{95.6} $\pm$ \textbf{0.0}\\
			
			\hline	
		\end{tabular}
	}			
	\caption{Accuracy Evaluation of Various Arabic NLP Models}
	\label{tab:acc}	
\end{table*}

For each dataset, we report the average results of five runs, each with different random seeds, to ensure statistical significance. The results for various Arabic NLP tasks are presented in Table \ref{tab:f1} and Table \ref{tab:acc} in terms of F1 and accuracy  metrics, respectively. From these results, we make the following observations:
\begin{enumerate}
\item 
%(1) 
Multilingual models such as mBERT and LaBSE are outperformed by Arabic-specific models that are pre-trained with larger vocabularies and more extensive language-specific datasets. This observation aligns with the findings of \citet{Abdul-MageedEN20}.
\item 
%(2)
Models that incorporate dialectal data during pre-training, such as MdBERT, CAMeL and MARBERT, not only excel in DID task, but also perform significantly across a broader range of Arabic NLP tasks. This suggests that the similarities among Arabic dialects may not always have positive effects on other tasks beyond ADI. The experimental results underscore the value of integrating more dialectal information during training, as the tokenizers in these models are likely to recognize more dialect-specific tokens, which are often unidentified in other models.

%Models like MdBERT, CAMeL and MARBERT, which integrate dialectal data during pre-training, excel not only in DID task but also across various Arabic NLP tasks. This underscores the significance of integrating more dialectal information during training to improve model performance.
\item 
%(3)
Despite being trained on less MSA text and fewer training steps, due to computational resource constraints, our model outperforms its alternatives in most tasks and achieves competitive performance in others. Although the improvements in tasks other than ADI are modest, they are significant given the inherent complexities of the Arabic language.
\end{enumerate}

In tasks beyond DID, AlcLaM may show modest improvements, but it introduces vital empirical factors like stability and statistical significance, supported by a t-test (\(p < 0.05\)). 
%MdBERT, CAMe, MARBERT, and our AlcLaM consistently demonstrate superior performance across a range of Arabic NLP tasks. These empirical findings clearly support our claim regarding the critical importance of incorporating Arabic dialectal data in the pre-training process.
MSA-Dialect PLMs consistently demonstrate superior performance across a range of Arabic NLP tasks. These empirical findings clearly support our claim regarding the critical importance of incorporating Arabic dialectal data in the pre-training process.
\section{Conclusion}

In this paper, we present AlcLaM, a novel BERT-based model trained specifically to address the challenge of Arabic dialectal variation. Leveraging a carefully curated corpus sourced from social media platforms, AlcLaM its alternatives across various Arabic NLP tasks, despite being trained on significantly less data. 
%This highlights the effectiveness of incorporating dialectal data in model training. 
For future work, expanding the dialectal vocabulary without increasing inference costs, inspired by Chinese character modeling.
%For future work, drawing inspiration from advancements in Chinese character modeling, expanding the Arabic dialectal vocabulary of existing PLMs without inflating inference costs could be explored. However, further investigation is required in this area.
%For future work, drawing inspiration from advancements in Chinese character modeling, expanding the Arabic dialectal vocabulary of existing PLMs without inflating inference costs could be explored. However, further investigation is required in this area.
%For future work, inspired from work in Chinese characters, expending the the Arabic dialectal vocabulary of existing PLMs without compromising the inference cost could be adopted; however, further investigation is need. 

\section*{Limitations}
Despite the advancements achieved by AlcLaM, it is important to acknowledge its current limitations:
\begin{itemize}
	\item AlcLaM is trained from scratch to build its vocabulary. However, incorporating weights of new dialectal vocabulary from existing Arabic PLMs and adjusting through continued training is a potential avenue for enhancement. Nevertheless, expanding the vocabulary size to encompass more dialectal tokens might lead to increased inference costs.
	\item Given that AlcLaM was trained on approximately 10\% of the training data used by its alternatives, due to computational resource constraints, its performance on generative tasks may not be as significant. Nonetheless, this limitation can be mitigated by continued training on our open-source AlcLaM model.
	
\end{itemize}

\bibliography{anthology,custom}
\appendix

\end{document}